OpenStudy (megannicole51):

Find the Taylor series about 0 for each of the functions below. Give the first three non-zero terms for (1+2x)^4

OpenStudy (megannicole51):

i tried expanding it out and making it 16x^4+32x^3+24x^2+8x+1 and taking the derivatives but it didnt work

ganeshie8 (ganeshie8):

two ways to do this :- 1) find derivatives using chain rule directly 2) use the known series (1+x)^n

ganeshie8 (ganeshie8):

\(f(x) = f(0) + x f'(0) + \frac{x^2}{2!}f^{''}(0) + ....\)

ganeshie8 (ganeshie8):

lets try the first method maybe.. find f', f"...

OpenStudy (megannicole51):

f'(x)=64x^3+96x^2+48x+8

OpenStudy (megannicole51):

f''(x)=192x^2+192x+48

ganeshie8 (ganeshie8):

hey u knw chain rule right ?

OpenStudy (megannicole51):

yes

OpenStudy (megannicole51):

4(1+2x)^3*2 right? or something like that

ganeshie8 (ganeshie8):

\(f(x) = (1+2x)^4\) \(f'(x) = 4(1+2x)^3*2 = 8(1+2x)^3\) \(f''(x) = 8*3(1+2x)^2*2 = 48(1+2x)^2\)

ganeshie8 (ganeshie8):

\(f(0) = (1+2*0)^4 = 1\) \(f'(0) = 8(1+2*0)^3 = 8\) \(f''(0) = 48(1+2*0)^2 = 48\)

OpenStudy (megannicole51):

okay!

ganeshie8 (ganeshie8):

plug them in the series

OpenStudy (megannicole51):

1+8x+48x^2/2! yay!!!

ganeshie8 (ganeshie8):

wat we're doing is the maclaurin series; hope u knw it... taylor series @ 0 becomes maclaurin series

OpenStudy (megannicole51):

can u help me with one more and then i am soooo going to bed! yeah my professor doesnt teach very well so I'm struggling this semester :/

OpenStudy (megannicole51):

finding the first 4 terms of sqrt(1-7x)

OpenStudy (megannicole51):

i know the first term is 1:)

ganeshie8 (ganeshie8):

great start :)

OpenStudy (megannicole51):

ahhh yeauuuh ima pro! lol

ganeshie8 (ganeshie8):

lets find f(0), f'(0), f"(0)... \(f(x) = \sqrt{1-7x}\) \(f'(x) = \frac{-7}{2\sqrt{1-7x}}\) \(f^{''}(x) = \frac{-49}{4(1-7x)^{3/2}}\) \(f^{'''}(x) = \frac{-1029}{8(1-7x)^{5/2}}\)

ganeshie8 (ganeshie8):

lets finish this first, there is a much easier way than this, but lets finish this first ok

ganeshie8 (ganeshie8):

\(f(0) = \sqrt{1-7*0} = 1\) \(f'(0) = \frac{-7}{2\sqrt{1-7*0}} = \frac{-7}{2}\) \(f^{''}(0) = \frac{-49}{4(1-7*0)^{3/2}} = \frac{-49}{4}\) \(f^{'''}(0) = \frac{-1029}{8(1-7*0)^{5/2}} = \frac{-1029}{8}\)

ganeshie8 (ganeshie8):

plug them in the series

ganeshie8 (ganeshie8):

\(f(x) = f(0) + x f'(0) + \frac{x^2}{2!}f^{''}(0) + ....\) \(f(x) = 1 -\frac{7}{2}x - \frac{49}{8}x^2 - \frac{1029}{48}x^3 - .... \)

OpenStudy (megannicole51):

yay i got it:)) thank you so much seriously uve helped so much!

ganeshie8 (ganeshie8):

np :) im sure u deserve some good sleep after so much hard work... gnite :)